(1) Just as a reminder, write out the following numbers: 1, 500 billion on z trillion Nellie sells 5 used bicycles. Their prices are: $150, $200, $250, $300, and $1100. What is the average (mean) price? The median price is the price of the middle item (the item at the half-way point). What is the median price? (2) (a) (b) The median is a better indicator of the middle of a group of items when there are a few items with a much higher or lower value than the rest (called outliers). In this case the higher-priced bicycle pulled the average (up/down) DECIMALS AND PERCENTAGES:
Answer #1

1. 1,500 billion = 1,500,000,000,000 = 1.5 trillion

Explanation: 1 billion = 1,000,000,000. So, 1,500 billion = 1,500 * 1,000,000,000 = 1,500,000,000,000.

Also, 1,000 billion = 1 trillion. So, 1,500 billion = 1.5 trillion.

2. Average price = $(150 + 200 + 250 + 300 + 1100)/5 = $2,000/5 = $400.

Median price is $250.

The high prices item put the average up.

Explanation: Average = sum of the items/no. of items.

Median is the value of the item in the medium position when the numbers are arranged in an ascending or descending order.

Very high values pull the average up and very low values pull the average down.